# An Improved Machine Learning Approach for RFI Mitigation in FAST-SETI Survey Archival Data

### v1

@@ 3.2 RFI Removal Results
Figure 4: Comparison result between the DBSCAN algorithm and the KNN algorithm for residual RFI mitigation. Data points marked in red represent signals identified and removed as residual RFI, while those in black are retained hits. The left panel shows the waterfall plot after applying the DBSCAN algorithm, demonstrating a residual RFI removal rate of 77.87%. The right panel displays the waterfall plot of the same dataset using the KNN algorithm, achieving a 70.43% removal rate (wang2023search). Notably, the green boxes in the left panel highlight RFI that DBSCAN effectively mitigation but KNN fails to identify, accounting for approximately 7.44%.
@@
3.3 Candidate Signal Selection
After mitigating the RFI, we select candidate ETI signals from the remaining 10508 hits. Characteristically, such candidate signals do not form large clusters in time-frequency scales; instead, they are narrowband in frequency and have a duration no longer than the telescope’s observation time for a single sky point during the drift scan. We also employ the DBSCAN algorithm to identify these small clusters, setting MinPts to 5 (much less than the threshold used to remove residual RFI), as follows the selection of MinPts for ETI candidate signal search by wang2023search and zhang2020first. With all simulated signals guaranteed to remain, the Eps is set to ${9.5}\times{10}^{-4}$ (Figure 6). After the candidate signal selection using DBSCAN algorithm, 364 clusters are identified, including 20 birdie clusters.
Following the Section 3.3 of wang2023search, the candidate clusters are further selected based on the following conditions:
@@ 3.3 Candidate Signal Selection
Through the above conditions, all 20 birdie clusters are marked which is consistent with the real number of birdie groups mentioned in Section 3.1, and 33 candidate clusters are selected, as shown in Figure 7. The final cluster count demonstrates a significant reduction compared to the 83 candidates reported by zhang2020first, and similar to (overlap substantially ) the 31 candidate clusters found by wang2023search. This result corroborates the our methodological validity.
To improve the accuracy of the screening results, we use the visual inspection (see wang2023search for details) to check whether the 33 candidate clusters above have obvious RFI characteristics. According to wang2023search and zhang2020first, since some selected clusters may actually be very close to other hits removed as RFI, these clusters need to be eliminated and regarded as parts of the RFI that were missed in previous steps. Specifically, wang2023search extracted and checked the raw data within distances of 0.1 MHz, 1000 s to the candidate groups (all previously identified RFIs are marked rather than removed in the raw data), and found 14 promising groups that do not seem to be part of large RFI clusters. Similarly, using this visual inspection method, we also select 14 interesting candidate groups (consistent with those identified by wang2023search) from the 33 candidate clusters, as shown in Figure 8. More detailed information of these candidate groups is shown in Table 2. These results show that the ETI signal is less likely to be misidentified as RFI after the removal of RFI using the DBSCAN algorithm due to the reduced amount of data. In addition, the method can not only remove more RFI, but also effectively improve the work efficiency of candidate signal selection.
@@
Figure 7: Selected ETI candidates (33, black dots) and identified birdies (20, red dots) in the frequency ($f$)-time ($t$) plane, with other filtered hits shown in grey. The methodology’s effectiveness is demonstrated by the recovery of all 20 “birdies” and a final candidate count of 33 (e.g., 83 by zhang2020first; 31 by wang2023search).
Figure 8: Visualizations of 14 interesting candidate groups. These candidates were initially identified based on ETI signal characteristics and subsequently passed visual inspection, a process that confirms they do not exhibit obvious RFI features. These 14 groups are part of 33 such candidates selected in this work and are equal to those previously found by wang2023search.
3.4 Candidate Signal Analysis
The 14 interesting candidate groups described above are mainly results obtained by machine learning algorithms, and further analysis still needs to be conducted. First, since the effective band of the FAST L-band receiver is 1050–1450 MHz (2011IJMPD..20..989N), we remove four candidate groups within the invalid bands (50 MHz wide each) at both ends of this range. Second, according to environmental monitoring of RFI at the FAST site, there are two sources of RFI in the observation band: civil aviation and navigation satellites (2021RAA....21...18W). There are five groups located in the civil aviation band (1030-1140 MHz) and one group located in the navigation satellite band (1176.45 $\pm$ 1.023 MHz, 1207.14 $\pm$ 2.046 MHz, 1227.6 $\pm$ 10 MHz, 1246.0-1256.5 MHz, 1268.52 $\pm$ 10.23 MHz, and 1381.05 $\pm$ 1.023 MHz). In addition, the frequency of Group No. 12 is located on the edge of one of the navigation satellite bands. Since satellite frequencies usually drift, we also consider this frequency to be approximately included in the navigation satellite band. Therefore, in order to eliminate the impact of this interference, we exclude seven candidate groups located within these frequency bands.
@@ 3.4 Candidate Signal Analysis
where $\delta$ is the declination. Since the declination of the signal source is approximately 0, $\cos\delta=1$. The HPBW decreases with the increase of frequency (2020RAA....20...64J). Within the effective frequency band, the lowest frequency corresponds to the maximum HPBW, which in turn corresponds to the longest duration. Taking the central beam (M01) as an example, when the frequency f=1060 MHz (the minimum frequency in Table 2 of 2020RAA....20...64J), the HPBW is $3.44^{\prime}$; in this case, $t_{max}=13.76s$. That is, the maximum duration of an ETI signal within a single beam is 13.76 s; if a signal persists longer than this duration, it should be identified as RFI. Therefore, we exclude four candidate groups in Table 2, as their durations are longer than the maximum value ($t_{max}$).
In conclusion, only Candidate Group No. 14 from Table 2 remain, and it cannot be ruled out as RFI based on current information. This candidate group has a short duration ($\sim 2.8$ s) and show no frequency drift. In fact, due to the Doppler effect, the frequency of a signal from an extraterrestrial origin will drift as the observation duration increases (tao2022sensitive). Therefore, we will conduct longer-duration observations to strive for a more comprehensive judgment in the future.
@@
4 Discussion
There are many clustering algorithms available, with HDBSCAN(Hierarchical Density-Based Spatial Clustering of Applications with Noise) (campello2013density) and OPTICS (Ordering Points to Identify the Clustering Structure) (ankerst1999optics) being notable variants and improvements of the DBSCAN algorithm. These advanced algorithms demonstrate effectiveness in handling variable density clusters and noisy data, with HDBSCAN, for example, successfully employed in recent RFI clustering studies(jacobson2025anomaly). We perform comparative tests to ensure our selection of DBSCAN for residual RFI removal is optimal for our specific ETI signal search pipeline.
@@ 4 Discussion
5 Conclusion
In this paper, we propose an improved machine learning approach (i.e. DBSCAN algorithm) for RFI mitigation based on the 5 hr data of FAST commissioning drift-scan survey in July 2019 (zhang2020first). Apart from mitigating the persistent and drifting narrowband RFI, we apply the DBSCAN algorithm to remove the residual RFI, which is usually challenging to mitigate. The results show that the DBSCAN algorithm can successfully identify and remove 36977 residual RFIs (accounting for $\sim$ 77.87%), whereas the KNN algorithm achieves only 70.43% removal rate (wang2023search). Under identical processing conditions and using the same dataset, DBSCAN also demonstrates a 24.85% reduction in execution time compared to KNN. Finally, through the candidate signal selection, we find 14 interesting candidate groups, consistent with the results of wang2023search. Following further analysis, one of these candidate groups is retained. These findings further verify the effectiveness of the DBSCAN algorithm.

### v2

@@ III.2 RFI Removal Results
Figure 4: Comparison result between the DBSCAN algorithm and the KNN algorithm for residual RFI mitigation. Data points marked in red represent signals identified and removed as residual RFI, while those in black are retained hits. The left panel shows the waterfall plot after applying the DBSCAN algorithm, demonstrating a residual RFI removal rate of 77.87%. The right panel displays the waterfall plot of the same dataset using the KNN algorithm, achieving a 70.43% removal rate (Wang et al., 2023). Notably, the green boxes in the left panel highlight RFI that DBSCAN effectively mitigation but KNN fails to identify, accounting for approximately 7.44%.
@@
III.3 Candidate Signal Selection
After mitigating the RFI, we select candidate ETI signals from the remaining 10508 hits. Characteristically, such candidate signals do not form large clusters in time-frequency scales; instead, they are narrowband in frequency and have a duration no longer than the telescope’s observation time for a single sky point during the drift scan. We also employ the DBSCAN algorithm to identify these small clusters, setting MinPts to 5 (much less than the threshold used to remove residual RFI), as follows the selection of MinPts for ETI candidate signal search by Wang et al. (2023) and Zhang et al. (2020). With all simulated signals guaranteed to remain, the Eps is set to ${9.5}\times{10}^{-4}$ (Figure 6). After the candidate signal selection using DBSCAN algorithm, 364 clusters are identified, including 20 birdie clusters.
Following the Section 3.3 of Wang et al. (2023), the candidate clusters are further selected based on the following conditions:
@@ III.3 Candidate Signal Selection
Through the above conditions, all 20 birdie clusters are marked which is consistent with the real number of birdie groups mentioned in Section III.1, and 33 candidate clusters are selected, as shown in Figure 7. The final cluster count demonstrates a significant reduction compared to the 83 candidates reported by Zhang et al. (2020), and similar to (overlap substantially ) the 31 candidate clusters found by Wang et al. (2023). This result corroborates the our methodological validity.
To improve the accuracy of the screening results, we use the visual inspection (see Wang et al. (2023) for details) to check whether the 33 candidate clusters above have obvious RFI characteristics. According to Wang et al. (2023) and Zhang et al. (2020), since some selected clusters may actually be very close to other hits removed as RFI, these clusters need to be eliminated and regarded as parts of the RFI that were missed in previous steps. Specifically, Wang et al. (2023) extracted and checked the raw data within distances of 0.1 MHz, 1000 s to the candidate groups (all previously identified RFIs are marked rather than removed in the raw data), and found 14 promising groups that do not seem to be part of large RFI clusters. Similarly, using this visual inspection method, we also select 14 interesting candidate groups (consistent with those identified by Wang et al. (2023)) from the 33 candidate clusters, as shown in Figure 8. More detailed information of these candidate groups is shown in Table 2. These results show that the ETI signal is less likely to be misidentified as RFI after the removal of RFI using the DBSCAN algorithm due to the reduced amount of data. In addition, the method can not only remove more RFI, but also effectively improve the work efficiency of candidate signal selection.
@@
Figure 7: Selected ETI candidates (33, black dots) and identified birdies (20, red dots) in the frequency ($f$)-time ($t$) plane, with other filtered hits shown in grey. The methodology’s effectiveness is demonstrated by the recovery of all 20 “birdies” and a final candidate count of 33 (e.g., 83 by Zhang et al. (2020); 31 by Wang et al. (2023)).
Figure 8: Visualizations of 14 interesting candidate groups. These candidates were initially identified based on ETI signal characteristics and subsequently passed visual inspection, a process that confirms they do not exhibit obvious RFI features. These 14 groups are part of 33 such candidates selected in this work and are equal to those previously found by Wang et al. (2023).
III.4 Candidate Signal Analysis
The 14 interesting candidate groups described above are mainly results obtained by machine learning algorithms, and further analysis still needs to be conducted. First, since the effective band of the FAST L-band receiver is 1050–1450 MHz (Nan et al., 2011), we remove four candidate groups within the invalid bands (50 MHz wide each) at both ends of this range. Second, according to environmental monitoring of RFI at the FAST site, there are two sources of RFI in the observation band: civil aviation and navigation satellites (Wang et al., 2021). There are five groups located in the civil aviation band (1030-1140 MHz) and one group located in the navigation satellite band (1176.45 $\pm$ 1.023 MHz, 1207.14 $\pm$ 2.046 MHz, 1227.6 $\pm$ 10 MHz, 1246.0-1256.5 MHz, 1268.52 $\pm$ 10.23 MHz, and 1381.05 $\pm$ 1.023 MHz). In addition, the frequency of Group No. 12 is located on the edge of one of the navigation satellite bands. Since satellite frequencies usually drift, we also consider this frequency to be approximately included in the navigation satellite band. Therefore, in order to eliminate the impact of this interference, we exclude seven candidate groups located within these frequency bands.
@@ III.4 Candidate Signal Analysis
where $\delta$ is the declination. Since the declination of the signal source is approximately 0, $\cos\delta=1$. The HPBW decreases with the increase of frequency (Jiang et al., 2020). Within the effective frequency band, the lowest frequency corresponds to the maximum HPBW, which in turn corresponds to the longest duration. Taking the central beam (M01) as an example, when the frequency f=1060 MHz (the minimum frequency in Table 2 of Jiang et al. (2020)), the HPBW is $3.44^{\prime}$; in this case, $t_{max}=13.76s$. That is, the maximum duration of an ETI signal within a single beam is 13.76 s; if a signal persists longer than this duration, it should be identified as RFI. Therefore, we exclude four candidate groups in Table 2, as their durations are longer than the maximum value ($t_{max}$).
In conclusion, only Candidate Group No. 14 from Table 2 remain, and it cannot be ruled out as RFI based on current information. This candidate group has a short duration ($\sim 2.8$ s) and show no frequency drift. In fact, due to the Doppler effect, the frequency of a signal from an extraterrestrial origin will drift as the observation duration increases (Tao et al., 2022). Therefore, we will conduct longer-duration observations to strive for a more comprehensive judgment in the future.
@@
IV Discussion
There are many clustering algorithms available, with HDBSCAN(Hierarchical Density-Based Spatial Clustering of Applications with Noise) (Campello et al., 2013) and OPTICS (Ordering Points to Identify the Clustering Structure) (Ankerst et al., 1999) being notable variants and improvements of the DBSCAN algorithm. These advanced algorithms demonstrate effectiveness in handling variable density clusters and noisy data, with HDBSCAN, for example, successfully employed in recent RFI clustering studies(Jacobson-Bell et al., 2025). We perform comparative tests to ensure our selection of DBSCAN for residual RFI removal is optimal for our specific ETI signal search pipeline.
@@ IV Discussion
V Conclusion
In this paper, we propose an improved machine learning approach (i.e. DBSCAN algorithm) for RFI mitigation based on the 5 hr data of FAST commissioning drift-scan survey in July 2019 (Zhang et al., 2020). Apart from mitigating the persistent and drifting narrowband RFI, we apply the DBSCAN algorithm to remove the residual RFI, which is usually challenging to mitigate. The results show that the DBSCAN algorithm can successfully identify and remove 36977 residual RFIs (accounting for $\sim$ 77.87%), whereas the KNN algorithm achieves only 70.43% removal rate (Wang et al., 2023). Under identical processing conditions and using the same dataset, DBSCAN also demonstrates a 24.85% reduction in execution time compared to KNN. Finally, through the candidate signal selection, we find 14 interesting candidate groups, consistent with the results of Wang et al. (2023). Following further analysis, one of these candidate groups is retained. These findings further verify the effectiveness of the DBSCAN algorithm.